# New Advances in Phonons: From Band Topology to Quasiparticle Chirality

### v1

@@ V.3.1 Weyl Phonons are Both Topological and Chiral
The left panels of Fig. 33 illustrate an example of a twofold-degenerate quadruple Weyl phonon in the material BaPtGe Zhang et al. (2020b); Li et al. (2021a); Zhang et al. (2025b). This phonon mode features four parallel helicoid surface states due to its Chern number of 4, showcasing the interplay between topology and chirality in such systems.
@@ V.3.2 Diagnosing Weyl Phonons by Pseudo-angular Momentum
Topological classifications for Weyl quasiparticles with (screw) rotational symmetry $\mathcal{C}_{n}$ have been established Fang et al. (2012); Tsirkin et al. (2017); Zhang et al. (2025a). The monopole charge of a Weyl phonon on high-symmetry lines or at high-symmetry points depends on the symmetry of the bands forming the Weyl phonon, specifically the PAM in $\mathcal{C}_{n}$-invariant systems, as well as the little point group where the Weyl point is located Zhang et al. (2025a). Consequently, PAM can be used to diagnose Weyl phonons in $\mathcal{C}_{n}$-invariant systems. In some previous studies, PAM has been used to identify chiral phonons defined by AM, as neither helicity nor AM is easily observable in experiments, yet there is no intrinsic relationship between PAM and AM (or helicity). Zhang et al. (2023b) pointed out that, in certain chiral systems, chiral phonons can be identified by detecting PAM in the circularly polarized Raman scattering, since Weyl phonons are chiral phonons in such systems.
@@
Figure 34 provides an example in $\mathcal{C}_{3}$-invariant systems, where $p$ represents the pure rotation part of PAM and $\chi$ denotes the Chern number (monopole charge) Tsirkin et al. (2017). The Weyl phonons at both $\Gamma$ and A, which are time-reversal invariant momenta, are double Weyl phonons with a monopole charge of $-2$, while those along the $\Gamma$-A path are conventional Weyl phonons with a monopole charge of $-1$. Thus, relationship between PAM and the Chern number for Weyl phonons can be established, providing a systematic framework for detecting chiral phonons in the circularly polarized Raman scattering process. Moreover, the selection rules for circularly polarized Raman scattering in systems with rotational symmetries $\mathcal{C}_{n}$ are illustrated in Fig. 35.
@@
Experimental observations of chiral phonons in $\mathcal{C}_{3}$ screw rotation systems have been demonstrated in both Te Zhang et al. (2023b) and $\alpha$-HgS Ishito et al. (2023b), supported by first-principles calculations and symmetry analysis. These studies provide concrete evidence of chiral phonons in materials with rotational symmetry, highlighting the interplay between symmetry, topology, and phonon circularly polarization. In the next subsection, we will use Te as an example to demonstrate that Weyl phonons are chiral phonons and explore how to observe such phonons using PAM.
@@
Figure 36: (a) and (b) are the different physical quantities for Te with the right-hand and left-hand structure, respectively. (a2) and (b2) shows the circular atomic motion for the same phonon modes but with different chiral crystal structures, showing opposite chiral motions in real space. (a3) and (b3) shows the pseudo-angular momentum for the same phonon modes with different chiral crystal structures. (a4) and (b4) shows the Chern number of the Weyl phonon in different chiral crystal structures. Thus, there is a close relationship between the chirality of crystal structure, the polarization of the atomic motion in the real space, the sign of pseudo-angular momentum, and the Chern number. Adapted from Ref. Zhang et al. (2025a).
V.3.3 Weyl Phonons are Typical Chiral Phonons: Te As an Example
@@
Figures 36 (a2) and (b2) depict the atomic motions for the lower branch of the Weyl phonon along the $\Gamma$-A direction in right- and left-handed crystals, respectively. Figures 36 (a3) and (b3) are the PAM, (a4) and (b4) are the Chern number of the Weyl phonon in Te with right- and left-handed structures. If the chirality of the crystal structure changes, the circularly polarization (obtained from the atomic motion), the sign of PAM and the Chern number will also be altered. Therefore, the chirality of the crystal structure is intrinsically linked to the sign of the AM (and thus the helicity), the PAM, and the Chern number, demonstrating the profound connection between structural chirality and the topological and chiral properties of phonons Zhang et al. (2023b, 2025b, 2025a); Bousquet et al. (2024).
The observation of topological phonons and chiral phonons will be reviewd in the next section.
VI Experimental Progresses of Topological and Chiral Phonons
While theoretical advances have significantly expanded the classification of topological phonons, parallel progress in understanding chiral phonons has also driven momentum in the field. In this section, we review selected experimental explorations to provide recent breakthroughs and challenges.
@@ VI.1.1 Topological Phonons Observed by IXS and INS
Soon after theoretical predictions of topological phonons in crystalline materials Zhang et al. (2018a), these states are experimentally observed by in FeSi Miao et al. (2018). As described in Section IV, FeSi has B-20 structure that breaks the inversion-symmetry. Its phonon excitations host spin-1 Weyl phonon at the $\Gamma$ point and charge-2 Dirac phonon at the $R$ point as shown in Fig. 23. These topological quasiparticles are experimentally determined using IXS. Figure 37 (a) shows the phonon dynamical structure factor of FeSi near the $R$ point. The second derivative plot of Fig. 37 (a) is shown in Fig. 37 (b). Green and red curves are DFT calculated phonon band dispersion. The overall excellent agreement between DFT calculation and IXS spectra supports the effective low-energy model for charge-2 Dirac phonon at the R point Zhang et al. (2018a). A detailed fitting of the IXS curve shown in Figs. 37 (c)-(f) further confirm the degenerate point and Dirac-like phonon band dispersion near the R-point. Following the same experimental protocol, the $\mathcal{PT}$-symmetry protected phononic nodal lines were observed in MoB2 Zhang et al. (2019b), where the IXS spectra have shown phonon band degenerate points that form an endless helical string along the $L$-direction.
@@
Figure 37: Evidence of charge-2 Dirac phonon in FeSi. (a), (b) Color plot and second derivative plot of phonon dynamical structure factor near the R point. (c) and (d) show high-statistic IXS spectra at $\bm{Q}$=(4.26, 0.26, 0.26) and (4.5,0.5,0.5), respectively. The dashed blue curves are the fitted result. (e) The charge-2 Dirac phonon in the 50 meV energy range. (f) shows the high statistic IXS spectra at $\bm{Q}^{\Gamma R}$, $\bm{Q}^{R}$, and $\bm{Q}^{RM}$. Adapted from Ref. Miao et al. (2018).
@@
For topological quasiparticles, including Weyl fermions Murakami (2007); Wan et al. (2011); Weng et al. (2015); Lv et al. (2015b); Xu et al. (2015); Pal (2011), phonons Zhang et al. (2018a, 2020b); Miao et al. (2018); Li et al. (2021a); Zhang et al. (2023b); Liu et al. (2020a); Xia et al. (2019); Ding et al. (2022); Liu et al. (2021e, 2019); Li et al. (2018), magnons McClarty (2022), and fractionalized excitations Wen (1995); Oshikawa and Senthil (2006); Senthil and Fisher (2001a); Song and Hermele (2015); Ghaemi et al. (2012); Levin and Stern (2012); Chen et al. (2015); Shirley et al. (2019); Senthil and Fisher (2001b); Levin and Stern (2009); Stern (2016); Sun and Mao (2021); Seradjeh et al. (2009); Vaezi (2013); Maciejko and Fiete (2015), the topological invariants are defined on bulk wave functions. This has motivated experimental studies beyond the phonon band structure. As shown in Eq. (18), the phononic wave function enters $S(\bm{Q},\omega)$ through the term $\mathbf{Q}\cdot\mathbf{e}^{i}_{\mathbf{q\sigma}}$. For a topological chiral wave function, the three components of $\mathbf{e}^{i}_{\mathbf{q\sigma}}$ are complex numbers with different phase factors, describing how atoms move out-of-phase with one another. Therefore, the chiral phonon wave function will results in a non-trivial interference of the real and imaginary component of $\bm{e}_{\bm{q\sigma}}^{i}$, which strongly modify the momentum dependent intensity distribution of $S(\bm{Q},\omega)$. One can visualize this effect by calculating $S^{\text{DFT}}(\bm{Q},\omega),\ R(\bm{Q},\omega)\ \text{and}\ I(\bm{Q},\omega)$ Li et al. (2021a). Here $R(\bm{Q},\omega)$ and $I(\bm{Q},\omega)$ consider only the real and imaginary part of $\mathbf{e}^{i}_{\mathbf{q\sigma}}$ in Eq. (18). The chiral phononic wave function can therefore be revealed by quantitatively comparing the experimental $S(\bm{Q},\omega)$ and $S^{\text{DFT}}(\bm{Q},\omega)$. This approach has been applied in the IXS study of BaPtGe that hosts twofold quadruple Weyl (TQW) phonons (see Fig.24) Li et al. (2021a). It has been shown that the wave functions near the twofold quadruple Weyl are primarily real along the $\text{X}_{1}-\Gamma-\text{X}_{2}$ and $\text{M}_{1}-\Gamma-\text{M}_{2}$ high-symmetry directions. In contrast, the phononic wave function along the $\text{R}_{1}-\Gamma-\text{R}_{2}$ direction is mainly imaginary Li et al. (2021a).
To understand the direction-dependent wave function near the twofold quadruple Weyl point at $\Gamma$ point, Li et al. (2021a) developed a simplified model that considers only four Pt atoms in the unit cell with locations $\bm{r}_{\text{Pt1}}=(c,c,c)$, $\bm{r}_{\text{Pt2}}=(-c+\frac{1}{2},-c,c+\frac{1}{2})$, $\bm{r}_{\text{Pt3}}=(-c,c+\frac{1}{2},-c+\frac{1}{2})$, $\bm{r}_{\text{Pt4}}=(c+\frac{1}{2},-c+\frac{1}{2},-c)$, where $c$ is an arbitrary value satisfying $c\leq 1$. Based on the group theory, the atomic motions at the $\Gamma$ point can be described as three non-degenerate basis states with irreducible representations of $\Gamma_{1}^{(1)}$, $\Gamma_{2}^{(1)}$, $\Gamma_{3}^{(1)}$, and three threefold degenerate basis states with irreducible representations of $\Gamma_{4}^{(3)}$. The basis states of $\Gamma_{2}^{(1)}$ and $\Gamma_{3}^{(1)}$ that form the TQW can be derived by imposing chiral cubic crystal symmetry and time-reversal symmetry:
@@
where $\omega{=e}^{2\pi i/3}$, $A$ is a real constant, and $B$ is a complex constant. The $q$-dependent wave functions can be obtained by diagonalizing Eq. (91) in the $\phi_{\Gamma_{2}}$ and $\phi_{\Gamma_{3}}$ basis states. Along the [100] and [011] directions, the diagonal components of Eq. (91) are zero, imposing purely real wave functions, $\psi_{1}={(e^{i\theta/2}\phi}_{\Gamma_{2}}+e^{-i\theta/2}\phi_{\Gamma_{3}})\ /\sqrt{2}$ and $\psi_{2}=-i{(e^{i\theta/2}\phi}_{\Gamma_{2}}-{e^{-i\theta/2}\phi}_{\Gamma_{3}})\ /\sqrt{2}$, where $\theta=\text{arg}(B)$. In contrast, along the [111] direction, the off-diagonal components are zero yielding $\psi_{1}=\phi_{\Gamma_{2}}$ and $\psi_{2}=\phi_{\Gamma_{3}}$, which have large imaginary components. This simplified model analysis thus explains the directional wave function reported in Li et al. (2021a).
Following the similar spirit, Jin et al. (2022) showed that by quantitatively analysing the phonon dynamical structure, the Chern number can also be extracted. As an example, Jin et al. (2022) considered a twofold Weyl Hamiltonian:
@@
where $\mathbf{V(G)}$ is a constant vector that does not sensitively depend on $\mathbf{q}$ but varies by changing $\mathbf{G}$. Equation 94 shows that the INS intensity is sensitive to the projection of $\mathbf{S(q)}$ on $\mathbf{V(G)}$. This effect has been confirmed by an INS study of MnSi Jin et al. (2022). As shown in Fig. 39, the averaged INS intensity over solid $\mathbf{q}$ spheres show maximum and minimum consistent with Eq. (94).

### v2

@@ V.3.1 Weyl Phonons are Both Topological and Circularly Polarized
The left panels of Fig. 31 illustrate an example of a twofold-degenerate quadruple Weyl phonon in the material BaPtGe zhang2020twofold ; li2021observation ; zhang2025weyl . This phonon mode features four parallel helicoid surface states due to its Chern number of 4, showcasing the interplay between topology and chirality in such systems.
@@
We highlight two prototypical examples that decouple topological nature from phonon polarization: (i) The $A_{1}$ phonon in tellurium at general $k$-points is non-topological yet circularly polarized, its AM arising from the chiral symmetry of the crystal structure. (ii) The nodal-line phonons in MoB2 are topological, characterized by a Berry phase and $\mathcal{PT}$-symmetry protection, but are linearly polarized with vanishing AM, serving to decouple these concepts.
@@ V.3.2 Diagnosing Weyl Phonons by Pseudo-angular Momentum
Topological classifications for Weyl quasiparticles with (screw) rotational symmetry $\mathcal{C}_{n}$ have been established fang2012multi ; tsirkin2017composite ; zhang2025thechirality . The monopole charge of a Weyl phonon on high-symmetry lines or at high-symmetry points depends on the symmetry of the bands forming the Weyl phonon, specifically the PAM in $\mathcal{C}_{n}$-invariant systems, as well as the little point group where the Weyl point is located zhang2025thechirality . Consequently, PAM can be used to diagnose Weyl phonons in $\mathcal{C}_{n}$-invariant systems. In some previous studies, PAM has been used to identify chiral phonons defined by AM, as neither helicity nor AM is easily observable in experiments, yet there is no intrinsic relationship between PAM and AM (or helicity). zhang2023weyl pointed out that, in certain chiral systems, chiral phonons can be identified by detecting PAM in the circularly polarized Raman scattering, since Weyl phonons are chiral phonons in such systems.
@@
Figure 32 provides an example in $\mathcal{C}_{3}$-invariant systems, where $p$ represents the pure rotation part of PAM and $\chi$ denotes the Chern number (monopole charge) tsirkin2017composite . The Weyl phonons at both $\Gamma$ and A, which are time-reversal invariant momenta, are double Weyl phonons with a monopole charge of 2-2, while those along the $\Gamma$-A path are conventional Weyl phonons with a monopole charge of 1-1. Thus, relationship between PAM and the Chern number for Weyl phonons can be established, providing a systematic framework for detecting circularly polarized phonons in the circularly polarized Raman scattering process. Moreover, the selection rules for circularly polarized Raman scattering in systems with rotational symmetries $\mathcal{C}_{n}$ are illustrated in Fig. 33 zhang2025thechirality ; watanabe2025symmetry .
@@
Experimental observations of circularly polarized phonons in $\mathcal{C}_{3}$ screw rotation systems have been demonstrated in both Te zhang2023weyl and $\alpha$-HgS ishito2023truly , supported by first-principles calculations and symmetry analysis. These studies provide concrete evidence of circularly polarized phonons in materials with rotational symmetry, highlighting the interplay between symmetry, topology, and phonon circularly polarization. In the next subsection, we will use Te as an example to demonstrate that Weyl phonons are circularly polarized phonons and explore how to observe such phonons using PAM.
@@
Figure 34: (a) and (b) are the different physical quantities for Te with the right-hand and left-hand structure, respectively. (a2) and (b2) shows the circular atomic motion for the same phonon modes but with different chiral crystal structures, showing opposite chiral motions in real space. (a3) and (b3) shows the pseudo-angular momentum for the same phonon modes with different chiral crystal structures. (a4) and (b4) shows the Chern number of the Weyl phonon in different chiral crystal structures. Thus, there is a close relationship between the chirality of crystal structure, the polarization of the atomic motion in the real space, the sign of pseudo-angular momentum, and the Chern number. Adapted from Ref. zhang2025thechirality .
V.3.3 Weyl Phonons are Typical Circularly Polarized Phonons: Te As an Example
@@
Figures 34 (a2) and (b2) depict the atomic motions for the lower branch of the Weyl phonon along the $\Gamma$-A direction in right- and left-handed crystals, respectively. Figures 34 (a3) and (b3) are the PAM, (a4) and (b4) are the Chern number of the Weyl phonon in Te with right- and left-handed structures. If the chirality of the crystal structure changes, the circularly polarization (obtained from the atomic motion), the sign of PAM and the Chern number will also be altered. Therefore, the chirality of the crystal structure is intrinsically linked to the sign of the AM (and thus the helicity), the PAM, and the Chern number, demonstrating the profound connection between structural chirality and the topological and chiral properties of phonons zhang2023weyl ; zhang2025weyl ; zhang2025thechirality ; bousquet2024structural .
The observation of topological phonons and circularly polarized phonons will be reviewd in the next section.
VI Experimental Progresses of Topological and Circularly Polarized Phonons
Theoretical advances have motivated experimental studies of topological and chiral phonons. In this section, we review main experimental progresses.
@@ VI.1 Topological Phonons
Similar to topological fermions, the experimental study of topological phonons have been focused on both band dispersion, eigenvalues $\omega_{\sigma}(\mathbf{q})$, and wave functions, eigenvectors $\mathbf{\epsilon}_{\mathbf{q}\sigma}$. As we described in Sec. I B, these physical quantities can be obtained by directly measuring the phonon dynamical structure factor using energy and momentum resolved scattering techniques.
@@ VI.1.1 Topological Phonons Observed by IXS and INS
Soon after theoretical predictions of topological phonons in crystalline materials zhang2018double , these states are experimentally observed in FeSi miao2018observation . As described in Section IV, FeSi has B-20 structure that breaks inversion-symmetry. Its phonon excitations host spin-1 Weyl phonon at the $\Gamma$ point and charge-2 Dirac phonons at the $R$ point as shown in Fig. 21. The dispersion of these topological quasiparticles are experimentally determined using IXS. Figure 35 (a) shows the IXS determined $S(\mathbf{Q},\omega)$ of FeSi near the $R$ point. The second derivative plot of Fig. 35 (a) is shown in Fig. 35 (b). Green and red curves are DFT calculated phonon band dispersion. The overall excellent agreement between DFT calculation and IXS spectra supports the effective low-energy model for charge-2 Dirac phonon at the R point zhang2018double . A detailed fitting of the IXS curve shown in Figs. 35 (c)-(f) further confirm the degenerate point and Dirac-like phonon band dispersion near the R-point. Following the similar experimental protocol, $\mathcal{PT}$-symmetry–protected phononic nodal lines were observed in MoB2 zhang2019phononic . The IXS spectra revealed phonon band degeneracy points forming a nodal ring along the $L$-direction of the reciprocal space. Because these nodal rings are protected by $\mathcal{PT}$ symmetry, they can take arbitrary shapes and extend continuously across multiple Brillouin zones.
@@
Figure 35: Evidence of charge-2 Dirac phonon in FeSi. (a), (b) Color plot and second derivative plot of phonon dynamical structure factor near the R point. (c) and (d) show high-statistic IXS spectra at $\bm{Q}$=(4.26, 0.26, 0.26) and (4.5,0.5,0.5), respectively. The dashed blue curves are the fitted result. (e) The charge-2 Dirac phonon in the 50 meV energy range. (f) shows the high statistic IXS spectra at $\bm{Q}^{\Gamma R}$, $\bm{Q}^{R}$, and $\bm{Q}^{RM}$. Adapted from Ref. miao2018observation .
@@
Topological quasiparticles, including Weyl fermions murakami2007phase ; wan2011topological ; weng2015weyl ; lv2015observation ; xu2015discovery ; pal2011dirac , phonons zhang2018double ; zhang2020twofold ; miao2018observation ; li2021observation ; zhang2023weyl ; liu2020symmetry ; xia2019symmetry ; ding2022charge ; liu2021charge ; liu2019ideal ; li2018coexistent , magnons mcclarty2022topological , and fractionalized excitations wen1995topological ; oshikawa2006fractionalization ; senthil2001fractionalization_PRL ; song2015space ; ghaemi2012fractional ; levin2012classification ; chen2015anomalous ; shirley2019fractional ; senthil2001fractionalization_PRB ; levin2009fractional ; stern2016fractional ; sun2021fractional ; seradjeh2009exciton ; vaezi2013fractional ; maciejko2015fractionalized , are defined on topological invariants that are derived from the bulk wave functions. This requires experimental measurements to go beyond the phonon band structure. As shown in Eq. (20), the phononic wave function enters $S(\bm{Q},\omega)$ through the term $\bm{Q}\cdot\bm{\epsilon}_{\mathbf{q\sigma}}(s)$. For a topological chiral wave function, the three components of $\bm{\epsilon}_{\mathbf{q\sigma}}(s)$ are complex numbers with different phase factors, describing how atoms move out-of-phase with one another. Therefore, the circularly polarized phonon wave function will give rise to non-trivial interferences of the real and imaginary component of $\bm{\epsilon}_{\mathbf{q\sigma}}(s)$, which strongly modify the momentum dependent intensity distribution of $S(\bm{Q},\omega)$. One can visualize this effect by calculating $S^{\text{DFT}}(\bm{Q},\omega),\ R(\bm{Q},\omega)\ \text{and}\ I(\bm{Q},\omega)$ li2021observation . Here $R(\bm{Q},\omega)$ and $I(\bm{Q},\omega)$ consider only the real and imaginary part of $\bm{\epsilon}_{\mathbf{q\sigma}}(s)$ in Eq. (20). The chiral phononic wave function can therefore be revealed by quantitatively comparing the experimental $S(\bm{Q},\omega)$ and $S^{\text{DFT}}(\bm{Q},\omega)$. This approach has been applied in the IXS study of BaPtGe that hosts twofold quadruple Weyl phonons described in Section IV-A1 li2021observation . It has been shown that the wave functions near the twofold quadruple Weyl are primarily real along the $\text{X}_{1}-\Gamma-\text{X}_{2}$ and $\text{M}_{1}-\Gamma-\text{M}_{2}$ high-symmetry directions. In contrast, the phononic wave function along the $\text{R}_{1}-\Gamma-\text{R}_{2}$ direction is mainly imaginary li2021observation . To understand the direction-dependent wave function near the twofold quadruple Weyl point at the $\Gamma$ point, li2021observation developed a simplified model that considers only four Pt atoms in the unit cell with locations $\bm{r}_{\text{Pt1}}=(c,c,c)$, $\bm{r}_{\text{Pt2}}=(-c+\frac{1}{2},-c,c+\frac{1}{2})$, $\bm{r}_{\text{Pt3}}=(-c,c+\frac{1}{2},-c+\frac{1}{2})$, $\bm{r}_{\text{Pt4}}=(c+\frac{1}{2},-c+\frac{1}{2},-c)$, where $c$ is an arbitrary value satisfying $c\leq 1$. Based on the group theory, the atomic motions at the $\Gamma$ point can be described as three non-degenerate basis states with irreducible representations of $\Gamma_{1}^{(1)}$, $\Gamma_{2}^{(1)}$, $\Gamma_{3}^{(1)}$, and three threefold degenerate basis states with irreducible representations of $\Gamma_{4}^{(3)}$. The basis states of $\Gamma_{2}^{(1)}$ and $\Gamma_{3}^{(1)}$ that form the TQW can be derived by imposing chiral cubic crystal symmetry and time-reversal symmetry:
@@
where $\omega{=e}^{2\pi i/3}$, $A$ is a real constant, and $B$ is a complex constant. The $q$-dependent wave functions can be obtained by diagonalizing Eq. (99) in the $\phi_{\Gamma_{2}}$ and $\phi_{\Gamma_{3}}$ basis states. Along the [100] and [011] directions, the diagonal components of Eq. (99) are zero, imposing purely real wave functions, $\psi_{1}={(e^{i\theta/2}\phi}_{\Gamma_{2}}+e^{-i\theta/2}\phi_{\Gamma_{3}})\ /\sqrt{2}$ and $\psi_{2}=-i{(e^{i\theta/2}\phi}_{\Gamma_{2}}-{e^{-i\theta/2}\phi}_{\Gamma_{3}})\ /\sqrt{2}$, where $\theta=\text{arg}(B)$. In contrast, along the [111] direction, the off-diagonal components are zero yielding $\psi_{1}=\phi_{\Gamma_{2}}$ and $\psi_{2}=\phi_{\Gamma_{3}}$, which have large imaginary components. This simplified model analysis thus explains the directional wave function reported in li2021observation .
Following the similar spirit, jin2022chern showed that by quantitatively analysing the phonon dynamical structure, the Chern number can also be extracted. As an example, jin2022chern considered a twofold Weyl Hamiltonian:
@@
where $\mathbf{V(G)}$ is a constant vector that does not sensitively depend on $\mathbf{q}$ but varies by changing $\mathbf{G}$. Equation 102 shows that the INS intensity is sensitive to the projection of $\mathbf{S(q)}$ on $\mathbf{V(G)}$. This effect has been confirmed by an INS study of MnSi jin2022chern . As shown in Fig. 37, the averaged INS intensity over solid $\mathbf{q}$ sphere show maximum and minimum intensity in agreement with Eq. (102).